Bibliography

[AW] Micahel Atiyah and C.T.C. Wall, Cohomology of groups in Algebraic number fields, Ian Cassels and A. Frohlich (ed.), pp. 94–115.

[BFH] D. Bump, S. Friedberg and J. Hoffstein, Nonvanishing theorems for L-functions of modular forms and their derivatives, Inventiones mathematicae, vol. 102 (1990), pp. 543–618.

[Car] Roger Carter, Finite groups of Lie type, John Wiley and Sons, New York, 1985.

[Cas] Ian Cassels, Global fields in Algebraic number fields, Ian Cassels and A. Frohlich (ed.), pp. 42–84.

[CF] Ian Cassels and A. Frohlich (ed.), Algebraic number fields, Academic Press, London, 1967.

[Co] Brian Conrad, personal correspondence, March 4, 1998.

[Co2] Brian Conrad, personal correspondence, March 5, 1998.

[CS] Gary Cornell and Joseph Silverman (ed.), Arithmetic geometry, Springer-Verlag, New York, 1986.

[CSS] Gary Cornell, Joseph Silverman and Glenn Stevens (ed.), Modular forms and Fermat’s last theorem, Springer-Verlag, New York, 1997.

[Cox] David Cox, Primes of the form x^{2} + ny^{2}, John Wiley and Sons, New York, 1989.

[DI] Fred Diamond and John Im, Modular forms and modular curves in Seminar on Fermat’s last theorem, V. Kumar Murty (ed.), pp. 39–113.

[FD] Benson Farb and R. Keith Dennis, Noncommutative algebra, Springer-Verlag, New York, 1993.

[Fr] A. Frohlich, Local fields in Algebraic number fields, Ian Cassels and A. Frohlich (ed.), pp. 1–41.

[GZ] Benedict Gross and Donald Zagier, Heegner points and derivatives of L-series, Inventiones mathematicae, vol. 84 (1986), pp. 225–320.

[Ha] Robin Hartshorne, Algebraic geometry, Springer-Verlag, New York, 1977.

[Ko] Victor Kolyvagin, Finiteness of E(ℚ) and � (E/ℚ) for a class of Weil curves, Math. of the USSR Izvestiya, vol. 32 (1989), pp. 523–542.

[La-Al] Serge Lang, Algebra, Addison-Wesley. [La-ANT] Serge Lang, Algebraic number theory, Springer-Verlag, New York, 1986. [La-EF] Serge Lang, Elliptic functions, Springer-Verlag, New York, 1987. [La-MF] Serge Lang, Introduction to modular forms, Springer-Verlag, Heidelberg, Germany, 1995.

[Ma] Hideyuki Matsumura, Commutative ring theory, Cambridge University Press, Cambridge, 1986.

[Maz] Barry Mazur, An introduction to the deformation theory of Galois representations in Modular forms and Fermat’s last theorem, Gary Cornell, Joseph Silverman and Glenn Stevens (ed.), pp. 243-311.

[Mi-AV] J.S. Milne, Abelian Varieties in Arithmetic Geometry, Gary Cornell and Joseph Silver-

man (ed.), pp. 103–150. [Mi-ADT] J.S. Milne, Arithmetic duality theorems, Academic Press, Boston, 1986. [Mi-EC] J.S. Milne, Étale cohomology, Princeton University Press, Princeton, New Jersey, 1980. [Mi-JV] J.S. Milne, Jacobian varieties in Arithmetic Geometry, Gary Cornell and Joseph Silver-

man (ed.), pp. 167–212.

[Mu] David Mumford, Abelian Varieties, Oxford University Press, Oxford, 1970.

[MM] M.R. Murty and V.K. Murty, Mean values of derivatives of modular L-series, Annals of Mathematics, vol. 133 (1991), pp. 447–475.

[Mur] V. Kumar Murty (ed), Seminar on Fermat’s last theorem, American Mathematical Society Press, Providence, 1995.

[PR] Vladimir Platonov and Andrei Rapinchuk, Algebraic groups and number theory, Academic Press, Boston, 1994.

[Pi] Richard Pierce, Associative algebras, Springer-Verlag, New York, 1982.

[Roh] David Rohrlich.

[Ro] Michael Rosen, Abelian varieties over ℂ in Arithmetic Geometry, Gary Cornell and Joseph Silverman (ed.), pp. 79–101.

[Ru-TS] Karl Rubin, Tate-Shafarevich groups and L-functions of elliptic curves with complex

multiplication, Inventiones mathematicae, vol. 89 (1987), pp. 527–560. [Ru-ES] Karl Rubin, Euler systems and Iwasawa theory. [Se-CA] Jean-Pierre Serre, A course in arithmetic, Springer-Verlag, New York, 1973. [Se-GC] Jean-Pierre Serre, Galois cohomology, Springer-Verlag, New York, 1997. [Se-LF] Jean-Pierre Serre, Local fields, Springer-Verlag, New York, 1979. [Se-EC] Jean-Pierre Serre, Propriétés galoisiennes des points d’ordre fini des courbes elliptiques,

Inventiones mathematicae, vol. 15 (1972), pp. 259–331.

[ST] Jean-Pierre Serre and John Tate, Good reduction of abelian varieties, Annals of Mathematics, vol. 68 (1968), pp. 492–517.

[Sh] Stephen Shatz, Group schemes, formal groups and p-divisible groups in Arithmetic geometry, Gary Cornell and Joseph Silverman (ed.), pp. 29–78.

[Shi] Goro Shimura, Introduction to the arithmetic theory of automorphic functions, Princeton University Press, Princeton, New Jersey, 1971.

[Si-AEC] Joseph Silverman, The arithmetic of elliptic curves, Springer-Verlag, New York, 1986. [Si-H] Joseph Silverman, The theory of height functions in Arithmetic geometry, Gary Cornell

and Joseph Silverman (ed.), pp. 151–166. [Si-2] Joseph Silverman, Advanced topics in the arithmetic of elliptic curves, Springer-Verlag,

New York, 1993.

[Ta] John Tate, Global class field theory in Algebraic number fields, J.W.S Cassels and A. Frohlich (ed.), pp. 162–203.

[Ti] Jacques Tilouine, Hecke algebras and the Gorenstein property in Modular forms and Fermat’s last theorem, Gary Cornell, Joseph Silverman and Glenn Stevens (ed.), pp. 327–342.

[Wei] Charles Weibel, An introduction to homological algebra, Cambridge University Press, Cambridge, 1994.

[Wes-Ide] Tom Weston, The idelic approach to number theory. [Wes-IR] Tom Weston, The inflation-restriction sequence: an introduction to spectral sequences.